OpenStudy (anonymous):

Use the information provided to write the standard form equation of each circle x^2 + y^2-20x+26y+260=0

OpenStudy (anonymous):

(x-h)^2+(y-k)^6 = r^2 Now rearrange your equation

OpenStudy (anonymous):

how did you get (y-k)^6?

OpenStudy (anonymous):

sorry it is square typo

OpenStudy (anonymous):

It's fine I was confused. So for this problem the answer would be (x-10)^2 + (y+13)^2=.... I don't know how to get the last r^2 part

OpenStudy (anonymous):

Yes this is correct now see you need to add and subtract 100 to get (x-10)^2 similarly need to add and subtract 169. so finally we will have 260-169-100

OpenStudy (anonymous):

Okay thanks! So the last r^2=-9?

OpenStudy (anonymous):

no r^2 =9 because you take it to the right side (x-10)^2 + (y+13)^2 = 9

OpenStudy (anonymous):

Oh!!! Thanks sooo much!

OpenStudy (anonymous):

np :)